OpenStudy (anonymous):

Simplify Square Root of 200x^2

OpenStudy (anonymous):

\[\sqrt[2]{200x^2}\] = \[\sqrt{200}\] * \[\sqrt{x^2}\] = \[\sqrt{25*4*2}\]*x = \[\sqrt{2}\]*5*2*x

OpenStudy (anonymous):

\[\sqrt{200x^2}\]Find Factors 200x^2 2 and 100x^2 2 and 50^2 2 and 25x^2 5 and 5 x and x The terms with "twins" can go outside the radicand \[2*5*x \sqrt{2}\] \[10x \sqrt{2}\]